In the circuit shown, assume the battery emf is 30
Answer #1

charging voltage across capacitor will be written as

Vc =Vs(1 - e^(-t/RC))

=> 18.5 = 30 (1 - e^(-t/10.512))

=> 1 - e^(-t/10.512) = 18.5/30 = 0.616

=> e^(-t/10.512) = 0.3833

=> t = -10.512*ln(0.3833) = 10.08 sec